OpenStudy (anonymous):

if the letter N stands for 5, tell the value N+7= N-=5 so 5+7=12? 2N is 25? 3N -2= 35-2=33?

OpenStudy (saifoo.khan):

o_O

OpenStudy (anonymous):

ill help...

OpenStudy (anonymous):

Do you get it? :D

OpenStudy (saifoo.khan):

Nope! :?

OpenStudy (anonymous):

ok well I found what N ='s and then I solved it o_o

OpenStudy (anonymous):

N=5 so 5+7=12

OpenStudy (anonymous):

hang on let me work it out :D

OpenStudy (anonymous):

:D Thanks lemme know if you are confused

OpenStudy (anonymous):

Do you know what 3N is ?

OpenStudy (anonymous):

dont you mean 2N?

OpenStudy (anonymous):

yea u put 3N do u mean 2 N

OpenStudy (saifoo.khan):

o_O

OpenStudy (anonymous):

Yeah sorry lol tofu Xd

OpenStudy (anonymous):

2N is what is says, if the space matters XD

OpenStudy (anonymous):

Yea i think you got it wright 33 is the anwser :D Yep 3n -2 = 35 -2 = 33 :D Good job

OpenStudy (anonymous):

Yay! I got all of em righto?

OpenStudy (anonymous):

Yep :D Good job :D

OpenStudy (anonymous):

YAYA!